# ImFace++: A Sophisticated Nonlinear 3D Morphable Face Model with Implicit Neural Representations

### v2

@@ 4.5 Ablation Study
On Disentangled Deformation Fields. To underscore the significance of the disentangled deformation learning process, we establish a baseline network that incorporates only one deformation field to universally learn face shape morphs. In this configuration, ${\mathbf{z}}_{exp}$ and ${\mathbf{z}}_{id}$ are concatenated and used as the input for the hyper net. Figure 12 (a) visually demonstrates the results. Despite some fine-grained details introduced by other designs, there is noticeable chaos in the reconstructed faces, especially for those with significant expressions. The quantitative results presented in Table III also indicate the significance of decoupled deformation learning.
@@
On Neural Blend-Field. We substitute the Neural Blend-Field in $\mathcal{E}$, $\mathcal{I}$, $\mathcal{T}$ with standard MLPs containing an equivalent number of parameters. These MLPs directly predict global deformations or SDF values for the entire face. As illustrated in Figure 12 (b), visible blurriness appears due to the limited capability to learn non-linear facial changes. The quantitative results in Table III further affirm the necessity of the Neural Blend-Field in achieving accurate representations.
@@
On Improved Embedding Learning. This strategy is introduced to facilitate the learning of more diverse and fine-grained facial expression deformations. As demonstrated in Figure 12 (c), when the number of expression embeddings is constrained to be the same as the expression categories, the generated expressions tend to be average. Furthermore, for exaggerated expressions, such as mouth stretching, the counterpart models struggle to converge to a reasonable state.
@@
On Refinement Displacement Field. RDF in ImFace++ is designed to learn and represent high-frequency details in facial shapes. A direct comparison between Figure 12 (d) and (e) illustrates that while ImFace can partially capture nonlinear facial deformations, it faces challenges in representing high-frequency details and personalized nuances. In contrast, ImFace++ excels at learning these intricate, subtle differences, by specially encoding them within the template space, exemplified by the sophisticated reproduction of features like the eyelid and wrinkles around the mouth. Furthermore, the quantitative validation presented in Table III underscores the crucial role of this newly introduced RDF module in ImFace++ for constructing a highly convincing 3D morphable model.
Figure 12: Qualitative ablation study results.
@@
Figure 13: Expression editing results. The faces present in the first row provide the source identity and facial detail embeddings.
@@ 4.6.1 Expression Editing
Similar to existing 3DMMs, ImFace++ serves as a comprehensive facial representation built upon prior distributions of facial expression and identity morphs, making it suitable for a variety of downstream applications. In Figure 13, we present expression editing results obtained from the FaceScape test set. The faces in the first row are the reconstructed ones providing source identities and neutral expressions, while the faces in the first column showcase the target expressions. In this figure, the first four rows depict reconstructed faces from the test set, and the last three rows feature generated faces from the train set. It is evident that facial expression editing can be easily accomplished by modifying the expression embeddings. These lifelike 3D faces demonstrate the impressive representational capabilities of ImFace++.
@@ 4.6.2 Facial Detail Transfer
In Fig. 14, ImFace++ demonstrates its capability to transfer high-frequency facial details across different identities and expressions. In particular, the target faces in columns 4 to 7 exhibit various eyelid types, a challenging task for explicit face representations. ImFace++ accomplishes the transfer of high-frequency details with impressive fidelity by merely swapping the detail embeddings. Furthermore, the model can also transfer facial wrinkles, such as frown lines and crow’s feet in the last two columns, to the source faces, showcasing its ability to capture and replicate fine-grained facial features with remarkable precision.
5 Conclusion
@@
This paper introduces ImFace++, a sophisticated nonlinear 3D morphable face model with INRs. ImFace++ excels in capturing complex facial shape variations through the use of two explicitly disentangled deformation fields associated with expression- and identity-variation, along with a refinement displacement field that encodes high-frequency facial geometry details, contributing to the creation of highly realistic facial representations. To enhance the precision in capturing detailed facial deformations and geometries, ImFace++ also introduces a Neural Blend-Field. Furthermore, an improved embedding learning strategy is proposed that enables more fine-grained expressions. Additionally, the paper presents an effective preprocessing pipeline that enables INRs to work with non-watertight facial surfaces for the first time. Experimental results demonstrate that ImFace++ is highly competent and outperforms the state-of-the-art methods in this domain.
Figure 14: Facial detail transfer results. ImFace++ is able to edit high-frequency facial details by modifying the corresponding embeddings. The faces in the first column provide the source identities and expression embeddings.
@@
While this work represents a significant advancement in 3D facial shape representation, the focus has primarily been on face geometry modeling, with less consideration given to facial appearance. Although a basic appearance model can be achieved by adding a color field, there leaves space for further exploration in developing a more comprehensive INR-based 3DMM that accurately represents facial appearance with realistic diffuse and specular albedo. Addressing this issue could be a significant direction for future research.
@@ Acknowledgments
This work is partly supported by the National Key R&D Program of China (2022ZD0161902), the National Natural Science Foundation of China (No. 62176012, 62202031, 62022011), Beijing Municipal Natural Science Foundation (No. 4222049), and the Fundamental Research Funds for the Central Universities.

### v3

@@ 4.5 Ablation Study
On Disentangled Deformation Fields. To underscore the significance of the disentangled deformation learning process, we establish a baseline network that incorporates only one deformation field to universally learn face shape morphs. In this configuration, ${\mathbf{z}}_{exp}$ and ${\mathbf{z}}_{id}$ are concatenated and used as the input for the hyper net. Fig. 13 (a) visually demonstrates the results. Despite some fine-grained details introduced by other designs, there is noticeable chaos in the reconstructed faces, especially for those with exaggerated expressions. The quantitative results presented in Tab. III also indicate the significance of decoupled deformation learning.
Figure. 13: Qualitative ablation study results.
@@
On Neural Blend-Field. We substitute the Neural Blend-Field in $\mathcal{E}$, $\mathcal{I}$, and $\mathcal{T}$ with standard MLPs containing an equivalent number of parameters. These MLPs directly predict global deformations or SDF values for the entire face. As illustrated in Fig. 13 (b), visible blurriness appears due to the limited capability to learn non-linear facial changes. The quantitative results in Tab. III further affirm the necessity of the Neural Blend-Field in achieving accurate representations.
@@
On Improved Embedding Learning. This strategy is introduced to facilitate the learning of more diverse and fine-grained facial expression deformations. As demonstrated in Fig. 13 (c), when the number of expression embeddings is constrained to be the same as the expression categories, the generated expressions tend to be average. Furthermore, for large expressions, such as mouth stretching, the counterpart models struggle to converge to a reasonable state.
@@
On Refinement Displacement Field. RDF in ImFace++ is designed to learn and represent high-frequency details in facial shapes. Direct comparison between Fig. 13 (d) and (i) illustrates that while ImFace can partially capture nonlinear facial deformations, it faces challenges in representing high-frequency details and personalized nuances. In contrast, ImFace++ excels at learning these intricate, subtle differences, by specially encoding them within the template space, exemplified by the sophisticated reproduction of features like the eyelid and wrinkles around the mouth. Furthermore, we conduct an experiment (w/o Corres.) to assess the impact of established correspondence on learning high-frequency details, where the displacement distance $\mathcal{D}:({\mathbf{p}_{b}},{\mathbf{z}}_{detail},l)\mapsto{\mathbf{d}}\in{\mathbb{R}}$ is modeled within the expression space instead of the aligned template space. The results are presented in Tab. III and Fig. 14, which clearly indicate the presence of artifacts around the mouth and eyes. This underscores the crucial role of the newly introduced RDF module and the impact of correspondence in ImFace++ for constructing a highly accurate 3D morphable model.
Figure. 14: Illustration of ablation study results on point-to-point correspondences. Zoom in for a better view.
On Hyper Net and Landmark Consistency Loss. We conduct an ablation study with two variations of the HyperNetwork architecture: 1) Concatenation (w/ Concat.): we concatenate the embeddings with coordinates and directly predict deformations and SDF using INRs. 2) Adaptive sentence normalization (w/ Ada.): We modify the vanilla implementation of adaptive sentence normalization by removing the normalization operation, enhancing the convergence with the sine activation function. As Tab. III shows, the Hyper Net architecture outperforms both variants, highlighting its efficacy in improving the model performance. Furthermore, we ablate the landmark consistency loss (w/o ${\cal L}_{lmk_{c}}$). The results in Tab. III indicate that it leads to a better accuracy of facial features and its absence incurs noticeable degradation, underscoring its importance.
@@
Figure. 15: Expression editing results. The faces present in the first row provide the source identity and facial detail embeddings.
@@ 4.6 Applications
ImFace++ leverages embeddings as input, aligning with the input requirements of traditional 3DMMs. This enables ImFace++ to effectively perform a broad range of downstream tasks traditionally handled by conventional 3DMM methods, such as facial registration and other related applications [106, 107, 108].
@@ 4.6.1 Expression Editing
Similar to existing 3DMMs, ImFace++ serves as a comprehensive facial representation built upon prior distributions of facial expression and identity morphs, making it suitable for a variety of downstream applications. In Fig. 15, we present expression editing results obtained from the FaceScape test set. The faces in the first row are the reconstructed ones providing source identities and neutral expressions, while the ones in the first column showcase the target expressions. In this figure, the first four rows depict reconstructed faces from the test set, and the last three rows feature the generated faces from the train set. It is evident that facial expression editing can be easily accomplished by modifying the expression embeddings. These lifelike 3D faces demonstrate the high representational capabilities of ImFace++.
Figure. 16: Facial detail transfer results. ImFace++ is able to edit high-frequency facial details by modifying the corresponding embeddings. The faces in the first column provide the source identity and expression embeddings.
@@ 4.6.2 Facial Detail Transfer
In Fig. 16, ImFace++ demonstrates its capability to transfer high-frequency facial details across different identities and expressions. In particular, the target faces in columns 2 to 4 exhibit various eyelid types, a challenging task for explicit face representations. ImFace++ accomplishes the transfer of high-frequency details with impressive fidelity by merely swapping the detail embeddings. Furthermore, the model can also transfer facial wrinkles, such as frown lines and crow’s feet in the last two columns, to the source faces, showcasing its ability to capture and replicate fine-grained facial features with a remarkable precision.
4.6.3 Texture Transfer
ImFace++ establishes high-quality correspondences, enabling effective texture transfer. As illustrated in Fig. 17, a UV texture map from the source face can be mapped to other target faces by utilizing the established correspondence between them, as described in Sec.4.3.2. This application further validates the robustness and accuracy of the correspondence quality in ImFace++.
4.6.4 Multi-view Reconstruction
ImFace++ is fully differentiable, enabling us to employ neural rendering techniques for reconstructing facial shapes from multi-view images. In particular, we use NeuFace [105] as our rendering framework, utilizing either ImFace or ImFace++ as the geometry module. Fig. 18 displays the recovered geometries. As can be seen, ImFace++ serves as a robust facial geometric prior model, enhancing ImFace by effectively capturing high-frequency features.
Figure. 17: Illustration of texture transfer. Leveraging the established correspondence, ImFace++ effectively transfers facial textures among different scans. Eyes are blurred for privacy.
Figure. 18: Multi-view reconstruction results. The Chamfer distance (CD) is listed at the right of each method.
Methods Forwarding Time(s)
i3DMM [33] 0.00552
NPHM [66] 0.00603
ImFace [87] 0.00456
ImFace++ 0.01961
TABLE IV: Forwarding time comparison. We evaluate the time required to forward 8,192 points using different INR-based models on a single NVIDIA Tesla A100 GPU. Neural Blend Fields are re-implemented in parallel following [111] to accelerate training and inference.
4.6.5 Ear-to-ear Face Reconstruction
We utilize the NPHM [66] dataset to assess the performance of our method on ear-to-ear facial shapes. In the Neural Blend Field, we set $k=7$ to refine facial regions and enhance the capacity of ImFace++ for representing facial geometries. We follow the procedure in Sec. 4.1 to sample points for training. As illustrated in Fig. 19, ImFace++ can be seamlessly extended to render ear-to-ear facial shapes, achieving high-quality results across various expressions and identities. Additionally, ImFace++ effectively handles high-frequency details, such as wrinkles and eyelids.
4.6.6 Texture Editing
ImFace++ captures meaningful geometry clues and high-quality correspondences, enabling effective texture editing. Specifically, we perform interpolation in the embeddings and UV maps, utilizing the point-to-point correspondence to map the texture onto each face. As demonstrated in Fig. 20, the texture exhibits significant changes in accordance with modifications in facial geometry, such as the appearance of wrinkles around the brow and mouth.
5 Discussion and Conclusion
@@
This paper introduces ImFace++, a sophisticated nonlinear 3D morphable face model with INRs. ImFace++ excels in capturing complex facial shape variations through the use of two explicitly disentangled deformation fields associated with expression- and identity-variations, along with a refinement displacement field that encodes high-frequency facial geometry details, contributing to the creation of highly realistic facial representations. To enhance the precision in capturing detailed facial deformations and geometries, ImFace++ also introduces a Neural Blend-Field. Furthermore, an improved embedding learning strategy is proposed that enables more fine-grained expressions. Additionally, the paper presents an effective preprocessing pipeline that enables INRs to work with non-watertight facial surfaces for the first time. Experimental results demonstrate that ImFace++ is highly competent and outperforms the state-of-the-art methods in this domain.
@@
While this work represents a significant advancement in 3D facial shape representation, the focus is primarily on face geometry modeling, with less consideration given to facial appearances. Although a basic appearance model can be achieved by adding a color field or using UV color maps, there leaves space for further exploration in developing a more comprehensive INR-based 3DMM that accurately represents facial appearances by integrating higher quality editable UV maps. Alternatively, exploring more sophisticated appearance models would also offer promising avenues to enhance the realism and detail of facial textures. Addressing this issue is a significant direction for future research. Moreover, as shown in Tab. IV, INR-based methods generally require more computational resources and time for training and inference, as they necessitate forwarding neural networks for each point to calculate the SDF values. However, researchers can implement fully-fused MLPs under the tiny-cuda-nn framework [109] or compute a cache of INRs [96] for faster training and inference. We leave this as our future work.
Figure. 19: Reconstruction results on ear-to-ear facial shapes (trained on the NPHM dataset).
Figure. 20: The illustration of texture editing results of ImFace++.
@@ Acknowledgments
This work is partly supported by the National Key R&D Program of China (2022ZD0161902), the National Natural Science Foundation of China (No. 62176012, 62202031, 62022011), Beijing Municipal Natural Science Foundation (No. 4222049), the Research Program of State Key Laboratory of Complex and Critical Software Environment, and the Fundamental Research Funds for the Central Universities.